OpenStudy (anonymous):

evaluate the discriminate and describe the nature of the roots for the equation

OpenStudy (anonymous):

??

OpenStudy (anonymous):

x^2+8x-10=0

OpenStudy (anonymous):

8^2 - 4*1*(-10) = 64 + 40 = 104 - a positive value means there are two unique real roots

OpenStudy (anonymous):

real and rational or real and irrational?

OpenStudy (anonymous):

they could be rational or irrational

OpenStudy (accessdenied):

The equation of the discriminant is: \( D = b^2 - 4ac \). Considering the discriminant within the quadratic formula: \[ x = \frac{-b \pm \sqrt{\color{green}D}}{2a} \] If \(D\) is positive, we have two cases: * If \(D\) is a perfect square, then \(\sqrt{D}\) is a positive integer, and we will have two rational solutions. * If \(D\) is not a perfect square, then \(\sqrt{D}\) is irrational, and we will have two irrational roots. If \(D\) is zero, then we have a single case: one rational root. If \(D\) is negative, then \(\sqrt{D}\) is not a real number, and we'll have two complex roots (no real roots).

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yes - sorry - in the above case the roots are real and irrational

OpenStudy (accessdenied):

You're welcome! :)